• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 11:21:52 on October30,2006

    Entry number 188553

    Followups:

    keyword=helicity 0 events dropped

    Helicity 0 events significantly dropped, while 1 and -1 increased.
    One difference from this run 1840 from the previous run was the increase of beam current from 50 uA to 80 uA, which should not affect helicity. Also before this run started, MCC did steering for Hall B.
    Asked MCC, they told me that Joe Grames did some work in the injector.
    See Elog 1345115. Joe did some work on Hall B IA and intensity.
    Joe used Hall A laser phase to adjust Hall B laser phase and then returned Hall A laser phase back to 144 deg. Maybe this is the reason
    for it?


    FIGURE 1